OpenStudy (anonymous):

Write an equation of a line that passes through the point (8, 4) and is parallel to the line y = 4x + 2. y = 4x − 28 y = 4x + 28 y = 1 over 4x − 2 y = 1 over 4x + 2 @mathmate

OpenStudy (mathmate):

@chrissy2430 Do you know what the slope is for the given equation y=4x+2?

OpenStudy (anonymous):

the slope is 4

OpenStudy (mathmate):

Excellent!

OpenStudy (mathmate):

Do you know the slope of a line which is parallel to this line?

OpenStudy (anonymous):

I forgot that part but I know the slope is 4.

OpenStudy (mathmate):

@chrissy2430 ?

OpenStudy (anonymous):

are u there? @mathmate

OpenStudy (mathmate):

The slope is indeed 4, you had a good guess. The rule is: lines that are parallel have the same slope. (remember this, so you don't have to guess again!)

OpenStudy (mathmate):

Now, does the line (with slope 4) pass through a given point? What is it?

OpenStudy (anonymous):

I think the answer is D? not to sure

OpenStudy (mathmate):

You need to know what you're doing before you can be sure of the answer, unless you show me how you got it. If not, I'll ask again: Now, does the line (with slope 4) pass through a given point? What is it?

OpenStudy (trojanpoem):

(x-x)/(y-y) = the slope

OpenStudy (anonymous):

I just told you I forgot how to o that and I need help... I know how to find the slope but forgot how to do the rest. can you please walk me threw it... im homeschooled and I really need help

OpenStudy (mathmate):

The answer to my question is in YOUR question, at the beginning of the post. Perhaps you did not read or understand my question.

OpenStudy (trojanpoem):

\[\frac{ x _{1} - x_{2} }{ y_{1}- y_{2} } = slope\]

OpenStudy (anonymous):

the slope formula is y2-yl1 over x2-x1 @TrojanPoem

OpenStudy (mathmate):

@TrojanPoem We already established the slope. If you don't mind, please let me finish the question, unless you want to take over.

OpenStudy (trojanpoem):

yep , that's why I hate the EQN :/

OpenStudy (mathmate):

I did not ask for the slope. YOUR question says the parallel line passes through a given point. I just want you to tell me which point it is.

OpenStudy (trojanpoem):

@mathmate Ops, my bad. I am out.

OpenStudy (anonymous):

idk that's why I said can you walk me threw it

OpenStudy (trojanpoem):

Bro, the line which is parallel to it got the same slope so it got the same equation

OpenStudy (mathmate):

Your question: "Write an equation of a line that passes through the point (8, 4) and is parallel to the line y = 4x + 2." MY question: "Now, does the parallel line (with slope 4) pass through a given point? What is it?"

OpenStudy (anonymous):

so would it be 4?

OpenStudy (mathmate):

I am asking for a point. Please re-read the question.

OpenStudy (trojanpoem):

It would be A

OpenStudy (mathmate):

To establish the parallel line, you already have the slope (=4). You need to know which point it passes through in order to find the equation of the line. Please re-read the original question (problem) to find out which point the parallel line should pass through.

OpenStudy (mathmate):

You need that point to find the equation of the line.

OpenStudy (anonymous):

A

OpenStudy (mathmate):

@chrissy2430 If you are only interested in answers, you will get stuck with the next problem. Can you confirm to me that you are not interested in knowing how to find the answer?

OpenStudy (trojanpoem):

He knows the answer as he said but he wanna check if he is right or not.

OpenStudy (mathmate):

@trojanpoem Where did he say he knows the answer? My bad if I didn't see it.

OpenStudy (anonymous):

all I was doing is seeing if I was doing it right and im not trying toust get the answer im trying to refresh everything I learned. I just became a homeschooler this year so everything I did in public i dont remember everything. all I asked was for help and if I said I didn't understand I don't see what the problem was for you to at la walk me threw it. but thank you f your help

OpenStudy (mathmate):

I cannot give you a monologue on how to solve it. I need you to get something out of it. Perhaps you'd be glad to work with @TrojanPoem , because he does not need feedback. @TrojanPoem Thank you for taking over!

OpenStudy (trojanpoem):

@mathmate I wish I am not destroying what you was trying to build , are I ?

OpenStudy (anonymous):

no your not lol your fine and thanks for everything @mathmate

OpenStudy (trojanpoem):

Post your questions , I will get out soon.

OpenStudy (mathmate):

@chrissy2430 What I wanted to tell you is when you have a line L with a given slope m that passes through a point \((x0,y0)\), the equation of the line is: \(y-y0 = m(x-x1)\) So for a line with slope 4 and passing through (8,4) would have an equation \(y-4=4(x-8)\) which simplifies to \(y-4=4x-32)\) and finally \(y=4x-28\) which is very easy to do. Since both you and @TrojanPoem wanted a fast answer, you got the answer without having been shown how to get it. Now you know, and I hope you take the time to understand it. There will always be similar questions in exercises, quizzes, and exams.